# A Weighted Hankel Approach and Cramér-Rao Bound Analysis for Quantitative Acoustic Microscopy Imaging

### v1

@@ 5.3 Ex vivo experiments
Experimental data were obtained using the 500 MHz QAM system described in [32] and collected from a 6-$\mu$m human lymph node section using a sampling frequency of 10 GHz. A 300-point RF signal was acquired at each spatial location using a 1-$\mu$m step size in both directions. An $60\!\times\!90$ $\mu$m spatial region of interest (ROI) was selected for analysis.
Figure 2: Left: Illustration of the normalized spectrum (3) for a bandwidth limited at -12 dB. Right: Magnitude in dB of the reference signal to illustrate -12 dB and -6 dB bandwidths.
@@ 5.4 Preprocessing
In [19], we investigated whether this preprocessing step is essential for the non-weighted HK-based method. While its use was important in simulations, its impact on real data was minimal. Nonetheless, we continue to apply the Cadzow filter to ensure fairness when comparing performance with the AR-based method, which requires this denoising step.
@@
The choice of bandwidth determines the range of frequencies where division in (3) is well-defined and affects the trade-off between the number of data points and the noise level in the truncated spectrum. The usable bandwidth typically includes frequencies within -6 to -12 dB of the center (or peak) frequency, depending on the sensitivity of the transducer and system technology, as obtained from $\mathcal{F}_{h_{0}}$ [1]. Fig. 2 illustrates the normalized spectrum from (3) with a bandwidth limited at -12 dB, alongside the magnitude in dB of the reference signal (left plot), highlighting the -12 dB and -6 dB bandwidth limits. Unless otherwise stated, a -12 dB bandwidth is used for analysis. To better understand its impact, we investigate the effects of bandwidth selection on both simulations and real data.
@@ 6.1 Simulations
Fig. 3 presents the estimation performance of the RHK (magenta solid line), HK (blue solid line), and AR (red solid line) estimators in simulated data, with variations in SNR, $Z$, $d$, $\alpha$, and the selected bandwidth (rows 1 to 5, respectively). The root-mean-square error (RMSE) of the estimated acoustic parameters is compared to the square root of the approximated CRB (black dashed line). Each column corresponds to the estimation of $c$, $Z$, $\alpha$, and $d$.
Figure 3: Simulations results. Columns represent RMSE results (calculated after outlier removal) for $c$, $Z$, $\alpha$ and $d$. Rows show different parameter-set variations for SNR, $Z$, $\alpha$ and $d$. RHK is in magenta, HK is in blue, AR is in red, and the square root of the CRB is in black dashed line.
Outliers (estimates falling outside the established range) were identified and excluded from the RMSE calculation in cases where any method failed. Fig. 4 depicts the percentage of outliers detected for each method in each scenario, with RHK represented by magenta bars, HK by blue bars, and AR by red bars.
In Fig. 3 (top row), it can be observed that as SNR increases, both RMSE and CRB decrease for all parameters, as expected. At high SNR values (above 60 dB), noise has minimal impact, allowing all methods to approach the CRB. However, at lower SNR, the gap between RMSE and CRB becomes more noticeable, with HK-based methods exhibiting lower RMSE values.
@@
Figure 4: Percentage of outliers detected for SNR, $Z$, $d$ and bandwidth variation. RHK is represented in magenta, HK in blue and AR in red.
@@
In Fig. 3 (third row), RMSE and CRB values decrease as $d$ increases. For all estimators, the RMSE approaches the CRB for large $d$. This confirms that thicker samples yield more accurate parameter estimates, while thinner samples (small $d$) present greater challenges for reliable estimation. Overall, all methods achieve similar RMSE values, except at $d\!=\!2.5\mu$m, where HK-based methods demonstrate lower RMSE.
@@
The fourth row of Fig. 3 demonstrates that as $\alpha$ increases, the RMSE slightly increases for the estimation of $c$ and $\alpha$. Notably, no outliers were detected in these experiments, thus the corresponding outliers report is omitted from Fig. 4.
The bottom row in Fig. 3 illustrates the impact of the bandwidth on the estimation performance. It is important to note that the CRB does not depend on the bandwidth, remaining constant across all scenarios. However, the gap between the RMSE and the CRB becomes more pronounced when smaller bandwidths are used. This can be explained due to a smaller bandwidth reduces the number of frequencies available for analysis, resulting in less data and, consequently, a more challenging estimation. In these cases, HK-based methods exhibited better performance than AR, while from a bandwidth of -10 dB, the performance of all estimators stabilizes and remains asymptotically close to the CRB.
@@
Figure 5: Outlier report in real data ROI. Left: Percentage of outliers detected across bandwidth variations, with RHK represented in magenta, HK in blue, and AR in red. Right: Box plot illustrating the distribution of the percentage of outliers for each method.
@@
To summarize, CRB curves play an important role, as expected, in predicting the behavior of the estimation algorithms and their dependence on parameter values and noise levels. Specifically, CRBc depends on all parameters, while CRBZ is influenced by $Z$ and $d$. Additionally, CRBα is dependent on all parameters except for $Z$, whereas CRBd depends on all parameters except for $\alpha$. Under favorable conditions, including high SNR, large $Z$ and $d$, and low $\alpha$, all methods show accurate performance with RMSE values approaching the CRB, demonstrating the reliability of the estimators. In challenging scenarios such as low SNR, small impedance contrasts, and thin samples, the proposed HK-based methods, particularly RHK, consistently outperform AR, achieving comparable or better estimation performance and significantly reducing the number of outliers.
6.2 Experiments
We conducted a study to evaluate the influence of bandwidth selection on the occurrence of unreliable estimates within the analyzed ROI. Fig. 5, on the left plot, presents the outlier ratios reported across different bandwidths, while on the right, box plots illustrate the distribution of the percentage of outliers for each method. All methods exhibited bandwidth dependency, with outlier ratios increasing as the bandwidth widened. This effect was more pronounced for the AR method compared to the HK methods. Interestingly, this behavior differs from the simulations, where a smaller number of data points was expected to degrade performance. In the case of real data, the discrepancy can be explained by the inclusion of frequencies farther from the central frequency when larger bandwidths are used. These frequencies introduce non-stationary and additional noise sources, such as electronic noise, which are not captured by the current model.
@@ 6.2 Experiments
Pixels with small (dark blue) or large (intense yellow) unreliable values are more prevalent in AR and HK than RHK, consistent with their larger outlier rates. Regions highlighted by red circles illustrate the improvements offered by RHK.
For all parameters, the median (red line) and interquartile range (blue box width) are similar across all algorithms. Notably, RHK consistently produces narrower distributions compared to HK and AR, reflecting improved stability and robustness. This results in smoother maps, with significant differences observed, particularly for the estimation of $\alpha$, where the RHK method demonstrates superior performance over its predecessor, HK.
@@ 7 Conclusions and Perspectives
In this work, we proposed a weighted HK-based framework for high-frequency signal processing in QAM. This method extends the baseline framework introduced in [19] by integrating a reweighting strategy based on redescending M-estimators. For the first time in the context of QAM, we derived CRBs, providing theoretical performance benchmarks for analyzing the accuracy of acoustic parameter estimation. Results in both simulated and experimental data at 500 MHz, demonstrated that the proposed approach consistently reduced outliers and improved parameter estimation accuracy compared to the standard AR approach, especially in challenging scenarios involving high noise levels or low impedance contrasts. These advancements come at a moderate computational cost, making the RHK framework a practical and effective alternative for real-world QAM applications.
Future work will seek to expand the scope and efficiency of the RHK framework. Testing datasets with varying noise levels and transducer frequencies, such as 1 GHz, will help identify scenarios where RHK surpasses HK, facilitating adaptive applications in more demanding conditions. Reducing the computational complexity of HK-based methods will be a key focus, with techniques like low-rank matrix factorization [34] offering promising solutions for scalability and real-time processing. Additional research will explore strategies to enhance robustness against variations in model order and bandwidth selection, as well as the integration of spatial information for joint or multivariate analyses to increase the flexibility of HK-based approaches. Finally, adopting deep unfolding methods [35] could improve optimization processes and robustness, paving the way for further advancements in this field.
@@ 9 Acknowledgments
We are grateful to the authors of [18] who have kindly provided us with the original code of their paper.

### v2

@@ 5.3 Ex vivo experiments
Experimental data were acquired using the 500 MHz QAM system described in [53]. Two biological specimens were analyzed: a 6 $\mu$m-thick section of human lymph node and a 6 $\mu$m-thick section of pig cornea. Data acquisition was performed at a sampling frequency of 10 GHz. At each spatial location, a 300-point RF signal was recorded using a lateral step size of 1 $\mu$m in both directions.
For analysis, regions of interest (ROIs) measuring $60\!\times\!90~{}\mu$m (lymph node) and $38\!\times\!99~{}\mu$m (pig cornea) were selected.
@@ 5.4 Preprocessing
In [38], we investigated whether this preprocessing step is essential for the non-weighted Hankel method. While its use was important in simulations, its impact on real data was minimal. Nonetheless, we continue to apply the Cadzow filter to ensure fairness when comparing performance with the AR-based method, which requires this denoising step.
@@
Figure 2: (a) Half-spectrum of the reference signal (substrate echo). (b) Half-spectrum of the QAM signal acquired from a sample region. (c) Normalized magnitude spectrum with -4 dB and -12 dB thresholds used for bandwidth estimation. (d) Ratio of QAM to reference spectra within the -4 dB bandwidth window.
The choice of bandwidth determines the range of frequencies where division in (4) is well-defined and affects the trade-off between the number of data points and the noise level in the truncated spectrum. The usable bandwidth typically includes frequencies within -6 to -12 dB of the center (or peak) frequency, depending on the sensitivity of the transducer and system technology, as obtained from $\mathcal{F}_{h_{0}}$ [2]. To illustrate the bandwidth selection, Fig. 2 shows example spectra used in our processing. The top plots correspond to the echo from the reference (no sample) and from a region with a sample. In the bottom row, we show how the bandwidth is defined using the -4 dB and -12 dB thresholds, and how the ratio is computed inside the selected frequency range. The impact of this choice is analyzed in both simulations and real measurements.
@@ 6.1 Simulations
We first evaluate the performance of the proposed methods using simulated data across varying SNR levels. Fig. 3 presents results for four estimators: proposed weighted Hankel (RHK, magenta), non-weighted Hankel [38] (HK, blue), autoregressive [25] (AR, red), and ESPRIT [37] (cyan).
Fig. 3 (a) shows the RMSE for each estimated parameter ($c$, $Z$, $\alpha$, and $d$) as a function of SNR. As expected, RMSE and the corresponding CRB both decrease with increasing SNR for all parameters. At high SNR values (above 60 dB), noise has little impact, allowing all methods to approach the theoretical limit given by the CRB. However, at lower SNRs, the discrepancy between RMSE and CRB becomes more pronounced, with slightly lower RMSE for Hankel methods and ESPRIT.
The percentage of outliers obtained for each method is reported in Fig. 3 (b). At low SNRs (e.g., 20-30 dB), all methods experience increased outlier rates, yet the Hankel methods demonstrate greater robustness. Specifically, at 20, 25, and 30 dB, RHK yields outlier rates of 84.5%, 66.5%, and 47%, respectively, while HK achieves 88.5%, 76.5%, and 49.5%. In contrast, ESPRIT and AR produce substantially higher outlier rates (ESPRIT: 99.0%, 95.5%, 66.5%; AR: 99.5%, 98.0%, 76.5%).
Figure 3: Estimation performance versus SNR using simulated 500 MHz QAM data. (a) RMSE for each estimated parameter ($c$, $Z$, $\alpha$, and $d$), after removal of outliers. The proposed RHK is shown in magenta, HK in blue, the AR in red, ESPRIT in cyan, and the square root of the CRB in black dashed line. (b) Percentage of outliers detected for each method versus SNR.
At 35-40 dB, RHK, HK and ESPRIT, have similar outlier rates close to 20% and 3.5%, respectively, while AR reported 27.0% and 6.0%.
Overall, ESPRIT and AR are more sensitive to noise and tend to produce a higher number of outliers under low SNR conditions. In contrast, Hankel approaches, particularly RHK, maintain better robustness, highlighting the advantage of the reweighting strategy in suppressing noise-induced estimation failures. Based on these findings, the remainder of our analysis focuses on comparing AR, as the established baseline for QAM.
Additional results at a fixed SNR of 50 dB are presented in Fig. 4 and Fig. 5, which examine RMSE behavior and outliers count as a function of tissue acoustic impedance $Z$, thickness $d$, and attenuation $\alpha$, as well as bandwidth selection.
Figure 4: Simulations results. Columns represent RMSE results (calculated after outlier removal) for $c$, $Z$, $\alpha$ and $d$. Rows show different parameter-set variations for $Z$, $\alpha$, $d$ and bandwidth. RHK is in magenta, HK is in blue, AR is in red, and the square root of the CRB is in black dashed line.
Figure 5: Percentage of outliers detected for $Z$, $d$ and bandwidth variation. RHK is represented in magenta, HK in blue and AR in red.
@@
In Fig. 4 (second row), RMSE and CRB values decrease as $d$ increases. For all estimators, the RMSE approaches the CRB for large $d$. This confirms that thicker samples yield more accurate parameter estimates, while thinner samples (small $d$) present greater challenges for reliable estimation. Overall, all methods achieve similar RMSE values, except at $d\!=\!2.5$ $\mu$m, where the proposed methods demonstrate lower RMSE.
@@
The third row of Fig. 4 demonstrates that as $\alpha$ increases, the RMSE slightly increases for the estimation of $c$ and $\alpha$. All methods yield similar RMSEs, with slightly higher errors observed for AR in some cases. In addition, a small gap between the RMSE and the CRB suggests that there is still room for improvement in estimation accuracy. Notably, no outliers were detected in these experiments, thus the corresponding outliers report is omitted from Fig. 5.
The bottom row in Fig. 4 illustrates the impact of the bandwidth on the estimation performance. It is important to note that the CRB does not depend on the bandwidth, remaining constant across all scenarios. However, the gap between the RMSE and the CRB becomes more pronounced when smaller bandwidths are used. This can be explained due to a smaller bandwidth reduces the number of frequencies available for analysis, resulting in less data and, consequently, a more challenging estimation. In these cases, Hankel methods exhibited better performance than AR, while from a bandwidth of -10 dB, the performance of all estimators stabilizes and remains asymptotically close to the CRB.
@@
To summarize, CRB curves play an important role, as expected, in predicting the behavior of the estimation algorithms and their dependence on parameter values and noise levels. Specifically, CRBc depends on all parameters, while CRBZ is influenced by $Z$ and $d$. Additionally, CRBα is dependent on all parameters except for $Z$, whereas CRBd depends on all parameters except for $\alpha$. Under favorable conditions, including high SNR, large $Z$ and $d$, and low $\alpha$, all methods show accurate performance with RMSE values approaching the CRB, demonstrating the reliability of the estimators. In challenging scenarios such as low SNR, small impedance contrasts, and thin samples, the proposed Hankel methods, particularly RHK, consistently outperform AR, achieving comparable or better estimation performance and significantly reduced number of outliers.
6.2 Experiments
We conducted an experiment to assess how bandwidth selection affects the occurrence of unreliable estimates within the lymph node ROI. In Fig. 6, the left plot shows the percentage of outliers detected for bandwidths ranging from -4 to -20 dB. The right plot presents boxplots summarizing the distribution of outlier rates for each method across this range.
Figure 6: Outlier analysis on the lymph node ROI. Left: Percentage of outliers detected across varying bandwidths (from -4 to -20 dB) for AR (red), HK (blue), and RHK (magenta). AR consistently yields higher outlier rates, particularly at wider bandwidths. Right: Boxplots showing the distribution of outlier percentages for each method. Reported $p$-values are from the Wilcoxon rank-sum test.
All methods exhibited sensitivity to bandwidth, with outlier rates generally increasing as the bandwidth expanded. To determine whether these differences were statistically significant, we performed pairwise comparisons using the non-parametric Wilcoxon rank-sum test [55]. This test was selected due to its robustness to non-normal distributions and its focus on median differences. The AR estimator showed significantly higher outlier rates compared to both HK and RHK, with $p$-values of 0.040 in both cases. No significant difference was observed between HK and RHK ($p=0.931$). These results suggest that Hankel methods are more robust than AR under experimental conditions.
Interestingly, this results contrasts with the simulation results, where reducing the number of frequency points (narrowing bandwidth) typically led to poorer performance. In real data, however, wider bandwidths include frequency components farther from the central frequency, which tend to introduce non-stationary behavior and additional noise, such as electronic interference, not accounted for in the current model.
@@ 6.2 Experiments
Figure 8: Distribution densities of estimated parameters ($c$, $Z$, and $\alpha$), including outliers, from the lymph node ROI. Each row corresponds to one parameter across the three methods (AR, HK, RHK).
Columns 2 to 4 show the corresponding estimation maps without outlier removal for AR, HK, and RHK methods. The rows represent $c$, $Z$ and $\alpha$, respectively. Pixels with small (dark blue) or large (intense yellow) unreliable values are more prevalent in AR and HK than RHK, consistent with their slightly larger outlier rates. Regions highlighted by red circles illustrate the visual improvements offered by RHK. While the visual differences may appear modest, this is primarily because the selected case represents a relatively favorable condition in which all methods, including AR, perform reasonably well. The advantages of the Hankel-based approaches are better demonstrated by their robustness in more challenging scenarios, as shown in Fig. 6. Nevertheless, we emphasize that even modest improvements are meaningful in the context of QAM, given the complexity of high-frequency ultrasound data and the inherent challenges of robust quantitative parameter estimation. Such advances support the broader goal of making QAM technology more accessible and practical.
As a complementary analysis, we present in Fig. 8 the true distribution densities of the estimated parameters, including outliers, for the lymph node ROI. To assess whether the estimation methods produce significantly different distributions, we conducted pairwise two-sample Kolmogorov-Smirnov tests [56], which are well suited for detecting differences between potentially multimodal distributions. The results, shown in Table 2, reveal that for attenuation, all comparisons between AR, HK, and RHK yielded statistically significant differences ($p<0.05$), indicating that the Hankel methods produce distributions distinct from AR. For speed of sound and impedance, however, no statistically significant differences were found.
Table 2: Results of two-sample Kolmogorov-Smirnov tests comparing the distributions of estimated parameters across methods. Statistically significant differences ($p<0.05$) are shown in bold.
Parameter Comparison p-value KS statistic
Attenuation $\alpha$ AR vs. HK $<$0.0001 0.0458
AR vs. RHK 0.0259 0.0290
HK vs. RHK $<$0.0001 0.0713
Speed of Sound $c$ AR vs. HK 0.1639 0.0220
AR vs. RHK 1.0000 0.0060
HK vs. RHK 0.1433 0.0226
Impedance $Z$ AR vs. HK 0.0694 0.0255
AR vs. RHK 0.9991 0.0073
HK vs. RHK 0.0638 0.0259
While the computational complexity of the frequency estimation algorithms was discussed in Section 3.2, we provide here the computational runtime for the experimental dataset as an illustrative example. On standard hardware (Intel i7, 32 GB RAM), processing a $60\times 90$ pixel region takes approximately 3 minutes for HK and 8 minutes for RHK, compared to under 1 minute for AR and ESPRIT. Although runtime is not a critical constraint in QAM applications, since processing is typically performed offline, accelerating QAM image formation remains an important open direction, though beyond the scope of this work. Potential avenues include adopting techniques developed for fast low-rank matrix completion in other imaging modalities, such as SVD-free solvers [57, 58], ADMM-free optimization methods [39, 59], and deep unfolding strategies [60].
@@
Similar results were observed in the selected pig cornea ROI. RHK yielded a slightly lower number of outliers across bandwidths compared to HK, while AR consistently reported higher outlier rates in all cases. Fig. 9 shows the acoustic parameter estimation maps for a bandwidth threshold of -4 dB, where the lowest outlier rates were obtained: AR = 12%, HK = 10%, and RHK = 9%. Some visual improvements can be observed with RHK, particularly in the estimation of $c$ and $Z$, compared to the other methods.
@@ 7 Conclusions and Perspectives
In this work, we proposed a weighted Hankel framework for high-frequency signal processing in QAM. This method extends our previous approach by integrating a reweighting strategy based on redescending M-estimators. For the first time in the context of QAM, we derived the Cramér-Rao bounds, providing theoretical performance benchmarks for analyzing the accuracy of acoustic parameter estimation. Results in both simulated and experimental data at 500 MHz, demonstrated that the proposed approach consistently reduced outliers and while keep or improve parameter estimation accuracy compared to the standard AR approach, especially in challenging scenarios involving high noise levels or low impedance contrasts. These advancements come at a moderate computational cost, making the RHK framework a practical and effective alternative for real-world QAM applications.
Future work will seek to extend the scope and efficiency of Hankel frameworks for QAM imaging. The approaches will be validated at high noise levels and different transducer frequencies, such as advanced 1 GHz QAM systems. We are also investigating strategies to accelerate the methods, with the goal of enabling faster QAM imaging. In addition, we plan to investigate joint estimation techniques that take into account spatial correlations in tissue, which may lead to more coherent and accurate parametric maps.
@@ 9 Acknowledgments
We are grateful to the authors of [28] who have kindly provided us with the original code of their paper. This work was supported by the National Institutes of Health (NIH) grant R01GM143388 (JM).